OpenStudy (horsegirl27):

Help I will medal and fan! I understand how to find the slope in a table. But how do I find the y-intercept? Also, I only slightly understand doing the same for equations written in separate forms. So explanations for that are wanted! Please help! And explain how to find them for an equation written in each form. Oh and how to write an equation from one form to another.

OpenStudy (horsegirl27):

\(\Huge\bf\color{red}{HELP~ME~PLEASE!!!}\)

OpenStudy (horsegirl27):

Lol I know it's a lot for one question but I really need help. So please explain how to do it!

OpenStudy (horsegirl27):

@sleepyjess please help me!!

OpenStudy (horsegirl27):

@iGreen @mathslover @Callisto @Conqueror @AnswerMyQuestions @triciaal @satellite73 @AidenS Please help me!

OpenStudy (horsegirl27):

*the table shows a function*

OpenStudy (sleepyjess):

Ok one thing at a time lol To find the y intercept I either graph it or just use the slope to find the pint when x = 0

OpenStudy (horsegirl27):

ok

OpenStudy (anonymous):

no repeated x

OpenStudy (horsegirl27):

?

OpenStudy (triciaal):

But how do I find the y-intercept? the y intercept is where the function crosses the y-axis at x = 0 so, set x = 0 find the value of y this is the y-intercept also b in the equation y = m x + b

OpenStudy (triciaal):

Oh and how to write an equation from one form to another. point slope form slope,m = y2 - y1/x2 - x1 for 2 points (x1, y1) and (x2, y2) the slope is the same for any 2 points on that line using the point (x, y) and one of the given points example (x1, y1) then m = y - y1/(x - x1) multiply each side of the equation by the LCD (x-x1) gives y - y1= m(x-x1) to put in slope intercept form expand the parenthesis and let y be the subject so y = mx -mx1 + y1 -mx1 + y1 = b y = m x + b

OpenStudy (horsegirl27):

Thanks so much!!!

OpenStudy (horsegirl27):

Any other explanations are welcome just to make sure I fully understand:)

OpenStudy (triciaal):

maybe you should get problems to practice with. math is just working problems. the concepts never change. same process with different numbers

OpenStudy (horsegirl27):

ok I get it now though thank you!!

OpenStudy (triciaal):

you are always welcome.

OpenStudy (horsegirl27):

:D